Two crates (91kg and 15 kg) are connected by a rope. The coefficient of kinetic force is 0.63. The crates have an acceleration of 1.7 m/s2. What is the force on the rope between the crates?
1
Expert's answer
2018-04-12T09:56:10-0400
The second Newton’s law:
Ma=F-T-μMg
ma=T-μmg
Thus,
T=m(a+μg)
T=15(1.7+0.63(9.8))=120 N